# Large Twist Angle Dependent Ultrafast Transient Dynamics of MoSe2/WSe2 van der Waals Heterostructures beyond the Mott Transition

### v2

@@ Introduction
Non-generate optical pump-optical probe measurements were performed using Ti: Sapphire ultrafast laser system (1 kHz repetition rate, 50 fs pulse width, M/s Spectra-Physics, Spitfire Ace) with the pump and probe energies of 3.1 eV and 1.55 eV, respectively. The pump and probe beams are passed collinearly through an objective lens with the NA = 0.9. The spot size of the beam at the sample point was measured to be 5 m. Experiments were carried out for different twist angles of MoSe2/WSe2 HS at two pump fluences of 2.0 and 2.8 mJ/cm2 at room temperature.
The MoSe2/WSe2 HS were prepared using the chemical vapor deposition method on a c-plane sapphire substrate via a two-step process similar to our earlier work [10]. The individual monolayers of MoSe2 and WSe2 have a triangular geometry as shown in the microscopic image of the sample (Figure 1a). The triangles with light (bright) contrast and bigger (smaller) in size represent WSe2 (MoSe2) monolayers as depicted by the black arrows in Figure 1a. A random alignment of triangular-shaped monolayers was obtained as a result of the CVD process, providing us with various twist angles. It has been verified that the TMDs with triangular geometry terminate with transition metal atoms at the edges using transmission electron microscopy[49] and scanning tunneling microscopy[50]. Therefore, the relative orientation of the layers provides us with the angle between the two layers as there is a direct correlation of crystal orientation with the triangular geometry of each monolayer[40, 49, 51]. The twist angle is evaluated as the angle between two lines collinear to the centroid and the vertex of the triangle[52]. One example of evaluating the twist angle is presented in Figure 1b, where red and blue outlined triangles represent MoSe2 and WSe2 monolayers, respectively. The lines collinear to the vertex and centroid are the angle bisector in the case of equilateral triangles, which gives the twist angle. The schemes for the front and side view of MoSe2/WSe2 HS for different twist angles are shown elsewhere [10].
@@
Figure S1 (supplementary information) shows the normalized intensities of M1, M2, and M3 with different twist angles. The black dashed lines represent the shift of the peak position of modes M1 or M2 at different twist angles. This systematic behavior of M1 and M2 can be seen in Figure 2b, where the dashed red lines (guides to the eyes) show the opposite slopes for M1 and M2 modes for four different ranges of the twist angles. We observed the softening of M1 and hardening of M2 mode for any non-zero twist angle as compared to the aligned geometry (depicted by blue circles for zero twist angle in Figure S1). Mode M2 starts stiffening as the twist angle is varied till 17∘, followed by softening till 32∘ and then stiffening again till 39∘ which is again followed by softening. In a counter behavior, the M1 mode softens as the twist angle is varied till 17∘, followed by stiffening till 32∘ and then softening till 39∘, which is again followed by stiffening. The softening and stiffening of the A1g mode (both M1 and M2) depend upon the carrier doping as well as the interlayer coupling [52, 56, 57]. In transition metal dichalcogenide MoS2, increasing the number of layers stiffens the A1g mode and softens the E2g mode due to increased dielectric screening of Coulomb interactions [58, 59]. It has been reported that the unidirectional transfer of electrons from monolayer WS2 to monolayer MoS2 leads to higher carrier density in MoS2 and lower carrier density in WS2, which can soften or stiffen the respective A1g modes[29]. For the similar staggered bands configuration of HS MoSe2/WSe2, unidirectional electron flow from WSe2 to MoSe2 increases (reduces) the carrier density in MoSe2 (WSe2) and leads to the softening (stiffening) of M1 (M2) modes. To understand the importance of twist angle, we have plotted the difference between the peak position of modes M1 and M2 with twist angles as shown in Figure 2c. A guide to the eye (solid red curve) shows a systematic variation of the shift of M1 with respect to M2 (or vice-versa) with maximum shift near the commensurate angles of 21.8∘ and 38.2∘. For perfectly aligned HS (=0∘), the electron transfer and the interlayer coupling lead to the softening of A1g of both M1 and M2 modes, if we compare these modes in monolayers of MoSe2 and WSe2 (Figure S1). However, with a non-zero twist angle, variation in the charge transfer and the interlayer coupling tunes the softening and the corresponding stiffening of M1 and M2 modes and shows an anomalous maximum shift near the commensurate angles. This can be easily understood as the maximum electron transfer and more dielectric screening of Coulomb interaction between electron and hole for the MoSe2/WSe2 HS at these specific angles, which are in the vicinity of commensurate angles. We also note the disappearance of M3 (2LA mode of WSe2) at the twist angles of 19∘ and 39∘ as depicted by small arrows in Figure S1 (panel iii), which needs further investigation.
@@
Figure 4 shows the type II band alignment in MoSe2/WSe2 HS depicting higher energy states of valence band maxima (VBM) and conduction band minima (CBM) of WSe2 with respect to that of MoSe2. The optical pump excitation using 400 nm (3.1 eV) laser pulse injects carriers into the conduction band much above the band gap for both constituent monolayers (MoSe2 Eg=1.57 eV, WSe2 Eg=1.67 eV) [60, 61], as shown in the left panel of Figure 4a. The energetic hot carriers undergo rapid electron-electron scattering and relax to the band edge, followed by a charge transfer between the MLs, creating both inter and intralayer excitons [34, 35]. However, at a photoexcitation density of 2 mJ/cm2 ( 6.11014 cm-2), these interlayer excitons undergo Mott transition, where they transform into a hot electron and hot hole plasma in MoSe2 and WSe2 monolayers, respectively [45, 46, 44, 43]. These carriers, localized to individual layers, emerge 1 ps after the photoexcitation, when most of the electrons (holes) get transferred to CBM (VBM) of MoSe2 (WSe2) [45, 35], as shown in Figure 4a. Localized interlayer charge carriers take longer to recombine compared to intralayer carriers, as the electrons and holes are confined to different layers rather than being in the same monolayer. The left panel of Figure 4b represents the radiative recombination of the intralayer electrons and holes in the individual monolayers, associated with (A1, ). For WSe2/MoSe2 HS with non-zero twist angle (42∘), observation of interlayer PL for excitation densities above Mott transition [45, 25] demonstrate the radiative recombination of interlayer electrons and holes, but only at a temperature of 4 K. However, Seyler et al. have reported that interlayer electrons and holes in MoSe2/WSe2 do not align at K valleys in the first Brillouin zone, emphasizing non-radiative recombination as a dominant process [27]. Therefore, we attribute the slow component of the relaxation dynamics (A2, ) to indirect radiative (phonon-mediated) and non-radiative recombination of interlayer electrons and holes, as illustrated in the middle panel of Figure 4b. Motivated by the suggestion of non-radiative Auger recombination of interlayer excitons at high excitation densities [62, 63], the A3 component is attributed to non-radiative interlayer Auger recombination. In this process, electrons and holes from different MLs recombine and transfer the energy to the third particle, such as an electron or a hole, as demonstrated in the right panel of Figure 4b.
@@
Figure 5a shows the twist-angle ()-dependent behavior of the intralayer photoexcited carriers (A1) and its recombination time () at a fluence of 2.0 mJ/cm2. A1 and exhibit no systematic dependence on twist angle as is expected for intralayer carriers recombination in individual monolayers. The recombination dynamics of the interlayer carriers is presented in Figure 5b, where amplitude A2 is twist-angle independent. The recombination time of interlayer localized electrons and holes, (panel ii of Figure 5b) exhibits minima near 20∘ and 38∘. The blue dots represent the data, while the solid blue line serves as a guide to the eye. The component associated with non-radiative Auger recombination of interlayer carriers, A3, also shows minima for 20∘ and 38∘ (Figure 5c). The trend of the fitting parameters at a fluence of 2.0 mJ/cm2 is similar at a fluence of 2.8 mJ/cm2, as shown in supplementary Figure S2.
@@
We offer a qualitative understanding of our results. We observed that is maximum at the twist angle of 30∘ (Figure 5b), i.e. the interlayer localized carriers take the longest time to recombine. This is due to the interlayer distance being maximum at the twist angle of 30∘ in TMDs HSs[28, 29]. Equally interesting is the observation of two minima of observed near 20∘ and 38∘. These twist angles are close to the commensurate angles of = 21.8∘ and 38.2∘ for MoSe2/WSe2 HS, where interlayer localized charge carriers or/and excitons can access an additional channel to recombine[27, 30]. Seyler et al. have demonstrated that for any non-zero twist angle between two layers, the holes in the valence band of WSe2 and electrons in the conduction band of MoSe2 at K points do not align in the first Brillouin zone. However, at the commensurate angle of 21.8∘, the valence band and conduction band align at K points in the second Brillouin zone, as shown in Figure 6, adapted from ref. [27]. At these specific angles, the interlayer charge carriers can recombine radiatively, and the momentum mismatch (due to non-zero twist angle) is compensated by the reciprocal lattice of two layers; the process known as Umklapp recombination [30, 27]. We attribute this Umklapp recombination in the vicinity of commensurate angles as a cause for the two minima observed for (Figure 5b). The two different fluences support the evidence of faster relaxation dynamics in the vicinity of commensurate twist angles. The additional radiative recombination channels for interlayer charge carriers result in lesser interlayer charge carriers involved in non-radiative Auger recombination, leading to a minimum in A3 near these commensurate angles (Figure 5c). To further investigate the minima for near the commensurate angles, the relaxation dynamics is studied as a function of excitation density for different twist angles.

### v3

# Twist Angle Dependent Ultrafast Transient Dynamics of MoSe2/WSe2 van der Waals Heterostructures beyond the Exciton Mott Transition

@@ Introduction
Non-generate optical pump-optical probe measurements were performed using Ti: Sapphire ultrafast laser system (1 kHz repetition rate, 50 fs pulse width, M/s Spectra-Physics, Spitfire Ace) with the pump and probe energies of 3.1 eV and 1.55 eV, respectively. The pump and probe beams are passed collinearly through an objective lens with the NA = 0.9. The spot size of the beam at the sample point was measured to be 5 m. The spot size is the same as the overlapping region of the heterostructure, and the resulting differential reflectivity () covers the entire overlapping region. Experiments were carried out for different twist angles of MoSe2/WSe2 HS at two pump fluences of 2.0 and 2.8 mJ/cm2 at room temperature. To investigate carrier relaxation dynamics above the exciton Mott transition, we use a probe beam with a fixed wavelength of 800 nm.
The MoSe2/WSe2 HS were prepared using the chemical vapor deposition method on a c-plane sapphire substrate via a two-step process similar to our earlier work [10]. The sapphire substrate were chosen because the strain effects are expected to be minimal (0.08) compared to silica substrate (0.38) [49]. The individual monolayers of MoSe2 and WSe2 have a triangular geometry as shown in the microscopic image of the sample (Figure 1a). The triangles with light (bright) contrast and bigger (smaller) in size represent WSe2 (MoSe2) monolayers as depicted by the black arrows in Figure 1a. A random alignment of triangular-shaped monolayers was obtained as a result of the CVD process, providing us with various twist angles. It has been verified that the TMDs with triangular geometry terminate with transition metal atoms at the edges using transmission electron microscopy[50] and scanning tunneling microscopy[51]. Therefore, the relative orientation of the layers provides us with the angle between the two layers as there is a direct correlation of crystal orientation with the triangular geometry of each monolayer[40, 50, 52]. The twist angle is evaluated as the angle between two lines collinear to the centroid and the vertex of the triangle[53]. One example of evaluating the twist angle is presented in Figure 1b, where red and blue outlined triangles represent MoSe2 and WSe2 monolayers, respectively. The lines collinear to the vertex and centroid are the angle bisector in the case of equilateral triangles, which gives the twist angle. The schemes for the front and side view of MoSe2/WSe2 HS for different twist angles are shown elsewhere [10].
@@
Figure S1 (supplementary information) shows the Raman spectra of M1, M2, and M3 modes with normalized intensities for different twist angles. The black dashed lines represent the shift of the peak position of modes M1 or M2 at different twist angles. We observed the softening of M1 and hardening of M2 mode for any non-zero twist angle as compared to the aligned geometry. Mode M2 starts stiffening as the twist angle is varied till 17∘, followed by softening till 32∘ and then stiffening again till 39∘ which is again followed by softening. In a counter behavior, the M1 mode softens as the twist angle is varied till 17∘, followed by stiffening till 32∘ and then softening till 39∘, which is again followed by stiffening. The softening and stiffening of the A1g mode (both M1 and M2) depend upon the carrier doping as well as the interlayer coupling [53, 57, 58]. In transition metal dichalcogenide MoS2, increasing the number of layers stiffens the A1g mode and softens the E2g mode due to increased dielectric screening of Coulomb interactions [59, 60]. It has been reported that the unidirectional transfer of electrons from monolayer WS2 to monolayer MoS2 leads to higher carrier density in MoS2 and lower carrier density in WS2, which can soften or stiffen the respective A1g modes[29]. For the similar staggered bands configuration of HS MoSe2/WSe2, unidirectional electron flow from WSe2 to MoSe2 increases (reduces) the carrier density in MoSe2 (WSe2) and leads to the softening (stiffening) of M1 (M2) modes. To understand the importance of twist angle, we have plotted the difference between the peak position of modes M1 and M2 with twist angles as shown in Figure 2b. Here, the solid blue line shows a systematic variation of the shift of M1 with respect to M2 (or vice-versa) with maximum shift near the commensurate angles of 21.8∘ and 38.2∘. For perfectly aligned HS (=0∘), the electron transfer and the interlayer coupling lead to the softening of A1g of both M1 and M2 modes, if we compare these modes in monolayers of MoSe2 and WSe2 (Figure S1). However, with a non-zero twist angle, variation in the charge transfer and the interlayer coupling tunes the softening and the corresponding stiffening of M1 and M2 modes and shows an anomalous maximum shift near the commensurate angles. This can be easily understood as the maximum electron transfer and more dielectric screening of Coulomb interaction between electron and hole for the MoSe2/WSe2 HS at these specific angles, which are in the vicinity of commensurate angles. We also note the disappearance of M3 (2LA mode of WSe2) at the twist angles of 19∘ and 39∘ as depicted by small arrows in Figure S1 (panel iii), which needs further investigation.
@@
Figure 4a shows the type II band alignment in MoSe2/WSe2 HS depicting higher energy states of valence band maxima (VBM) and conduction band minima (CBM) of WSe2 with respect to that of MoSe2. The optical pump excitation using 400 nm (3.1 eV) laser pulse injects carriers into the conduction band much above the band gap for both constituent monolayers (MoSe2 Eg=1.57 eV, WSe2 Eg=1.67 eV) [61, 62], as shown in the top left panel of Figure 4a. The energetic hot carriers undergo rapid electron-electron scattering and relax to the band edge, followed by a charge transfer between the MLs, creating both inter and intralayer excitons [34, 35]. However, at a photoexcitation density of 2 mJ/cm2 ( 6.11014 cm-2), these interlayer excitons undergo Mott transition, where they transform into a hot electron and hot hole plasma in MoSe2 and WSe2 monolayers, respectively [45, 46, 44, 43]. These carriers, localized to individual layers, emerge 1 ps after the photoexcitation, when most of the electrons (holes) get transferred to CBM (VBM) of MoSe2 (WSe2) [45, 35], as shown in the top right panel of Figure 4a. Localized interlayer charge carriers take longer to recombine compared to intralayer carriers, as the electrons and holes are confined to different layers rather than being in the same monolayer. The bottom left panel of Figure 4a represents the radiative recombination of the intralayer electrons and holes in the individual monolayers, associated with (A1, ). For WSe2/MoSe2 HS with non-zero twist angle (42∘), observation of interlayer PL for excitation densities above Mott transition [45, 25] demonstrate the radiative recombination of interlayer electrons and holes, but only at a temperature of 4 K. However, Seyler et al. have reported that interlayer electrons and holes in MoSe2/WSe2 do not align at K valleys in the first Brillouin zone, emphasizing non-radiative recombination as a dominant process [27]. Therefore, we attribute the slow component of the relaxation dynamics (A2, ) to indirect radiative (phonon-mediated) and non-radiative recombination of interlayer electrons and holes, as illustrated in the lower middle panel of Figure 4a. Motivated by the suggestion of non-radiative Auger recombination of interlayer excitons at high excitation densities [63, 64], the A3 component is attributed to non-radiative interlayer Auger recombination. In this process, electrons and holes from different MLs recombine and transfer the energy to the third particle, such as an electron or a hole, as demonstrated in the lower right panel of Figure 4a.
@@
Figure 5a shows the twist-angle ()-dependent behavior of the intralayer photoexcited carriers (A1) and its recombination time () at a fluence of 2.0 mJ/cm2. A1 and exhibit no systematic dependence on twist angle as is expected for intralayer carriers recombination in individual monolayers. The recombination dynamics of the interlayer carriers is presented in Figure 5b, where amplitude A2 is twist-angle independent. The recombination time of interlayer localized electrons and holes, (panel ii of Figure 5b) exhibits minima near 20∘ and 38∘. The blue dots represent the data, while the solid blue line is a cubic spline. The component associated with non-radiative Auger recombination of interlayer carriers, A3, also shows minima for 20∘ and 38∘ (represented by the solid blue line in Figure 5c). The trend of the fitting parameters at a fluence of 2.0 mJ/cm2 is similar at a fluence of 2.8 mJ/cm2, as shown in supplementary Figure S2.
@@
We offer a qualitative understanding of our results. We observed that is maximum at the twist angle of 30∘ (Figure 5b), i.e. the interlayer localized carriers take the longest time to recombine. This is due to the interlayer distance being maximum at the twist angle of 30∘ in TMDs HSs[28, 29]. Equally interesting is the observation of two minima of observed near 20∘ and 38∘. These twist angles are close to the commensurate angles of = 21.8∘ and 38.2∘ for MoSe2/WSe2 HS, where interlayer localized charge carriers or/and excitons can access an additional channel to recombine[27, 30]. Seyler et al. have demonstrated that for any non-zero twist angle between two layers, the holes in the valence band of WSe2 and electrons in the conduction band of MoSe2 at K points do not align in the first Brillouin zone. However, at the commensurate angle of 21.8∘, the valence band and conduction band align at K points in the second Brillouin zone, as shown in Figure 4b, adapted from ref. [27]. At these specific angles, the interlayer charge carriers can recombine radiatively, and the momentum mismatch (due to non-zero twist angle) is compensated by the reciprocal lattice of two layers; the process known as Umklapp recombination [30, 27]. We attribute this Umklapp recombination in the vicinity of commensurate angles as a cause for the two minima observed for (Figure 5b). The two different fluences support the evidence of faster relaxation dynamics in the vicinity of commensurate twist angles. The additional radiative recombination channels for interlayer charge carriers result in lesser interlayer charge carriers involved in non-radiative Auger recombination, leading to a trend that shows minima in A3 near these commensurate angles (Figure 5c). To further investigate the minima for near the commensurate angles, the relaxation dynamics is studied as a function of excitation density for different twist angles.